# SSFO: Self-Supervised Faithfulness Optimization for Retrieval-Augmented Generation

### v1

@@ 3.5 Encouraging Benign Probability Displacement with SSFO-$\lambda$
Empirical Validation of $\lambda$’s Effect. In Fig.˜4, we investigate the impact of varying $\lambda$ from 1.0 to 1.5 across multiple context-based question-answering benchmarks: NQ-Swap, NQ-Open, MemoTrap, and ELI5. As $\lambda$ increases we observe a consistent improvement in performance across all evaluated tasks. Pearson correlation coefficients $r$ reveal a positive relationship between $\lambda$ and performance on all datasets. For instance, span EM on NQ-Swap rises by about 1.6 points (from 77.2% to 78.8%), MemoTrap gains 1.7 points (from 74.8% to 76.5%). These results confirm that strategically amplifying the weight on the ungrounded (parametric) response via $\lambda>1$ indeed encourages a benign likelihood displacement and yields more faithful outputs.
Gradient Analysis. To further understand how this modification encourages the desired displacement, we analyze the gradient of this loss function. Following the derivation in dposhift , the gradient with respect to parameters $\theta$ is:
$\nabla_{\theta}\mathcal{L}_{\text{SSFO}-\lambda}=-\mathbb{E}\Big{[}c^{\prime}_{1}\big{(}\underbrace{\nabla_{\theta}\log\pi_{\theta}(y_{c}^{\prime}|x,c)}_{\text{increase likelihood of }y_{c}^{\prime}}-\underbrace{\lambda\nabla_{\theta}\log\pi_{\theta}(y_{p}|x,c)}_{\text{decrease likelihood of }y_{p}}\big{)}\Big{]},$ (5)
where $c^{\prime}_{1}$ is a positive coefficient. We present a detailed derivation of Eq.˜5 in Section˜C.1. Compared to the standard DPO update, Eq.˜4 applies a stronger negative weight ($\lambda$ where $\lambda>1$) to the gradient component associated with the parametric response $y_{p}$. Therefore, this parameter leads to a more pronounced suppression of the likelihood of the parametric response y during optimization.
@@ 4 Experiments
Datasets: To comprehensively evaluate faithfulness, we assess model performance across several dimensions.
Metrics: For short-form QA datasets (NQ-Open, NQ-Swap, MemoTrap, SQuAD), we adopt a standard zero-shot setting simulating a RAG scenario where the model answers queries based on the provided context without in-context learning’s demonstration. Performance is measured using the span Extraction Matching (span EM) score (a prediction is deemed correct if any segment of the generated output precisely matches one of the reference answers). For the long-form generation dataset ELI5, we report ROUGE scores rouge comparing the generated response against the ground-truth answer to assess content quality. For instruction following (FollowBench), we report Consistent Satisfaction Levels (CSL) followbench , which measures how many consecutive levels of instruction hardness a model can satisfy.
Models and Baselines: To ensure the generality of our approach, we conduct experiments using three families of open-source large language models LLaMA 3 touvron2023llama , Qwen 2.5 qwen2 and, Mistral as base models. We compare the proposed method against the strong publicly available methods focused on improving faithfulness: DECORE gema2024decore , ChatQA liu2025chatqa , Trust-Align trust_align , and Context-DPO context-dpo .
To construct the preference dataset, we randomly sample 900 query–context pairs from the MS MARCO datasetMS_MARCO , following the procedure outlined in Section˜3.2. Full implementation details are provided in Section˜D.4.
@@ 4.1 Faithfulness Evaluation Results
Instruct Model Method Implement
@@
NQ-Swap Memo-Trap NQ-Open SQuAD Eli5
Span EM $\uparrow$ Span EM $\uparrow$ Span EM $\uparrow$ Span EM $\uparrow$ R-1 F1 $\uparrow$ R-2 F1 $\uparrow$ R-L F1 $\uparrow$
Llama-3-8B Base model \ \ 73.54% 73.60% 80.15% 88.20% 26.88% 9.12% 23.88%
Decoding-Stratagy DECORE 80.53% 74.40% 82.03% 84.90% 28.45% 10.96% 25.65%
Post-training ChatQA 67.70% 30.60% 76.80% 88.50% 28.61% 12.67% 26.26%
Trust-Align 75.56% 70.95% 77.38% 50.90% 11.36% 10.55% 20.15%
Context-DPO 82.76% 72.90% 82.86% 89.90% 28.15% 10.43% 24.98%
SSFO 81.23% 76.28% 84.40% 89.00% 30.40% 11.80%   27.01%
SSFO-$\lambda$  ✗ 82.81% 78.38% 85.69% 90.90% 31.35% 13.37%   28.37%
Qwen2.5-7B Base model \ \ 79.35% 54.19% 82.29% 90.30% 23.08% 6.06% 20.55%
Decoding-Stratagy DECORE 81.93% 54.56% 83.76% 82.80% 24.31% 7.01% 21.46%
Post-training Trust-Align 79.69% 53.71% 77.93% 80.30% 23.15% 6.45% 20.24%
Context-DPO 82.13% 55.34% 83.13% 91.80% 23.68% 6.53% 21.13%
SSFO 84.18% 57.66% 83.88% 92.00% 24.63% 7.17%   21.83%
SSFO-$\lambda$ 84.88% 60.77% 84.48% 93.30% 28.53% 10.40%   25.66%
Mistral-7B Base model \ \ 67.76% 34.34% 79.13% 84.80% 23.20% 5.74% 20.38%
Decoding-Stratagy DECORE 78.17% 30.68% 86.52% 85.30% 23.90% 6.48% 21.19%
Post-training Context-DPO 79.62% 33.20% 80.68% 86.50% 24.97% 7.02% 22.00%
SSFO 86.66% 37.22% 87.53% 89.00% 32.67% 13.47%   28.35%
SSFO-$\lambda$ 85.48% 46.91% 90.32% 88.50% 39.22% 21.36%   37.24%
SSFO and SSFO-$\lambda$ achieve state-of-the-art faithfulness across multiple datasets and models, as shown in Table˜1:
  • Both variants of Self-Supervised Direct Preference Optimization (SSFO and SSFO-$\lambda$) substantially improve contextual faithfulness over the base models. For example, SSFO shows an improvement of 5-20% regarding Robustness over parametric knowledge, highlighting the effectiveness of self-supervised faithfulness optimization regarding response quality. We report further evaluation results across models with varying parameter sizes in Appendix˜E.
@@
  • SSFO-$\lambda$ attains state-of-the-art performance compared to both post-training-based baselines (ChatQA, Trust-Align, Context-DPO) and decoding strategy-based baseline (DECORE). These results confirm that explicitly encouraging benign likelihood displacement not only enhances faithfulness but also improves overall response quality.
[…]
Crose-language Response Quality Instruction Following
XQuAD(ES) DuReader(CN) FollowBench
Instruct Model Method Implement
Training Data
Required
Span EM $\uparrow$ Span EM $\uparrow$ CSL $\uparrow$
Base model \ \ 78.60% 78.80% 2.54
Decoding-Stratagy DECORE \ 83.87% 79.89% 2.46
ChatQA $\sim$30k 77.98% 72.05% 1.04
Trust-Align $\sim$15k 20.17% 8.12% 0.12
Context-DPO $\sim$5k 83.03% 84.40% 2.46
SSFO $\sim$800 83.10% 84.90%   2.70
Llama-3-8B Post-training SSFO-$\lambda$ $\sim$800 84.12% 83.56%   2.50
Base model \ \ 78.90% 81.50% 2.68
Decoding-Stratagy DECORE \ 80.08% 76.57% 2.56
Trust-Align $\sim$15k 75.21% 73.61% 0.58
Context-DPO $\sim$5k 79.92% 82.78% 2.64
SSFO $\sim$800 79.83% 83.27%   2.70
Qwen2.5-7B Post-training SSFO-$\lambda$ $\sim$800 81.76% 87.72%   2.62
@@
Table 3: Faithfulness improvement case study: comparing the base model, Context-DPO and SSFO on equivalent English–Spanish (ES) query–answer pairs. This example from XQuAD artetxe2020cross demonstrates that SSFO improves faithfulness over the base model and preserves faithfulness cross-lingually. Blue texts highlight context-grounded information; Red texts mark unfaithful hallucinations.
@@
Base Model (EN) Cultural Imperialism is often referred to as cultural hegemony or cultural dominance. Hallucination
Context-DPO (EN) Cultural Imperialism is often referred to as the spread of a dominant culture’s influence. This is also known as the use of soft power.                               Partially Context-Faithful
@@ 4.2 Generalization Across Tasks and Languages
SSFO can enhance multi-language faithfulness. We evaluate the generalization ability of SSFO in Table˜2, and results show it can achieve cross-lingual faithfulness using only an English-based training set. For instance, on Llama-3-8B, SSFO increases Span EM scores by 6.10% on DuReader (Chinese) and 5.52% on XQuAD (Spanish) compared to the base model. In contrast, heavily supervised methods like ChatQA and Trust-Align exhibit decreased performance on these non-English QA datasets. This shows that by training the model to prioritize context knowledge over parametric knowledge, it learns a principle of contextual adherence that can transfer across languages.
SSFO minimally impacts instruction following capability. Requiring only a few hundred self-supervised data examples, SSFO largely preserves, and even slightly enhances, the model’s instruction following capabilities.The CSL scores on FollowBench indicate that models fine-tuned with SSFO retain comparable general instruction-following ability to the original base instruct models. We provide several cases that SSFO retains strong general instruction following ability under context-based scenarios in Table˜4. In contrast, other post-training approaches such as Trust-Align improve faithfulness at the cost of degrading general generative abilities (e.g., CSL dropping to 0.12 for LLaMA-3-8B).
Table 4: Case study from FollowBench followbench : showing that SSFO retains strong instruction-following capabilities under a complex, context-based NLP task.
Instruction Analyze the provided text to pinpoint a category and the sentiment (positive, neutral, or negative) it emanates. Additionally, perform named entity recognition to emphasize notable entities and also identify the core topic discussed. Select the category from this array: company, educational institution, artist, athlete, office holder, means of transportation, building, natural place, village, animal, plant, album, film, or written work.
@@
Trust Align I apologize, but I couldn’t find an answer to your question in the search results.
@@ 4.3 Data Efficiency Analysis
Figure 5: Data efficiency study: SSFO requires about 60% of data (400–500 examples) to achieve 85% of the total possible performance gain over the base model.
To measure how many self-supervised preference examples SSFO actually needs, we subsample the preference dataset in 10% increments. As shown in Fig.˜5, we evaluate the average performance gain (an average improvement over the base instructed model). Both SSFO models cross the 85% performance threshold by approximately 50–60 % of the data ( 400–500 examples). We hypothesize that this efficiency stems from SSFO’s ability to leverage data generated by the model itself. This self-generated data inherently matches the model’s response style, thereby eliminating stylistic discrepancies and allowing the model to concentrate on optimizing the faithfulness gap. We compare the training examples from SSFO with other post-training methods in Section˜D.3.
5 Limitation
@@ 6 Conclusion
This work addressed the critical challenge of faithfulness hallucination in Retrieval-Augmented Generation (RAG) systems, where existing methods often introduce significant computational overhead or rely on costly external supervision. We introduced Self-supervised Faithfulness Optimization (SSFO), a novel and efficient self-supervised alignment method. SSFO uniquely leverages the LLM’s own outputs—contrasting responses generated with external contexts against those relying solely on parametric knowledge—to create preference pairs for alignment. Our experiments across diverse benchmarks show that SSFO significantly enhances model faithfulness and robustness against parametric knowledge, achieving state-of-the-art performance compared to existing methods. Furthermore, SSFO exhibits strong generalization capabilities, improving faithfulness even in cross-lingual settings using only English training data, while crucially preserving the model’s general instruction-following abilities.
Appendix A Related Work
A.1 Faithfulness Hallucination of Large Language Models
Hallucination in LLMs can be generally categorized into two types: factuality hallucination, where generated content deviates from established world knowledge (e.g., claiming “Mars has oceans”), and faithfulness hallucination, where the generated response is inconsistent with the provided context(e.g., misrepresenting a source document’s information). [9]

### v2

@@ 3.3.2 Encouraging Benign Probability Displacement with SSFO-$\lambda$
Empirical Validation of $\lambda$’s Effect. As shown in Fig.˜4, we investigate the impact of varying $\lambda$ from 1.0 to 1.5 across multiple context-based question-answering benchmarks: NQ-Swap, NQ-Open, MemoTrap, and ELI5. As $\lambda$ increases, we observe a consistent improvement in performance across all evaluated tasks. Pearson correlation coefficients $r$ reveal a positive relationship between $\lambda$ and performance on all datasets. For instance, span EM score on MemoTrap rises by 2.1 points (from 76.2% to 78.3%); NQ-Swap gains 1.2 points (from 81.2% to 82.5%). These results confirm that strategically amplifying the weight on the ungrounded (parametric) response via $\lambda>1$ (to encourage benign likelihood displacement) indeed yields a more faithful response.
[…]
$\nabla_{\theta}\mathcal{L}_{\text{SSFO}-\lambda}=-\mathbb{E}\Big[c^{\prime}_{1}\big(\nabla_{\theta}\log\pi_{\theta}(y_{c}^{\prime}|x,c)-\underbrace{\lambda\nabla_{\theta}\log\pi_{\theta}(y_{p}|x,c)}_{\text{decrease likelihood of }y_{p}}\big)\Big],$ (5)
where $c^{\prime}_{1}$ is a positive coefficient. We present a detailed derivation of Eq.˜5 in Section˜C.1. Compared to the standard DPO update, Eq.˜4 applies a stronger negative weight ($\lambda$ where $\lambda>1$) to the gradient component associated with the parametric response $y_{p}$. Therefore, this parameter leads to a more pronounced suppression of the likelihood of the parametric response during optimization.
@@ 4 Experiments
Datasets: To comprehensively evaluate faithfulness, we assess model performance across several dimensions. (1) For evaluating Robustness against conflicting parametric knowledge,we follow prior work Gema et al. (2024); Shi et al. (2024), using MemoTrap Liu & Liu (2023) and NQ-Swap Longpre et al. (2021).(2) For Response Quality, we evaluate on the context-based short-form QA datasets NQ-Open Lee et al. (2019) and SQuAD Rajpurkar et al. (2016), as well as the long-form generation datasets ELI5 Fan et al. (2019) and WikiPassageQA Cohen et al. (2018). (3) To assess the generalization ability of the proposed methods, we benchmark Cross-language Response Quality using DuReader He et al. (2018) and XQuAD Artetxe et al. (2020), and Instruction Following Ability using FollowBench Jiang et al. (2024).
Metrics: For short-form QA datasets (NQ-Open, NQ-Swap, MemoTrap, SQuAD), we adopt a standard zero-shot setting simulating a RAG scenario where the model answers queries based on the provided context. Performance is measured using the span Extraction Matching (span EM) score (a prediction is deemed correct if any segment of the generated output precisely matches one of the reference answers). For the long-form generation dataset ELI5, we report ROUGE scores Lin (2004) to quantify lexical overlap between the generated responses and the reference answers. We also report the LLM-Faithfulness Score (LFS). The LFS is calculated using GPT-4 Achiam et al. (2023) to classify outputs as faithful, partially faithful, or unfaithful (see the prompt in Table˜7), and the score is defined as the number of faithful generations divided by the total number of generations. For instruction following (FollowBench), we report Consistent Satisfaction Levels (CSL) Jiang et al. (2024), which measures how many consecutive levels of instruction hardness a model can satisfy.
Models and Baselines: To ensure the generality of our approach, we conduct experiments using three families of open-source large language models: LLaMA 3 Instruct Touvron et al. (2023), Qwen 2.5 Instruct Yang et al. (2024), and Mistral Instruct. We compare the proposed method against the strong methods focused on improving faithfulness: CAD Shi et al. (2024), DECORE Gema et al. (2024), ChatQA Liu et al. (2025), Trust‑Align Song et al. (2025), Context‑DPO Bi et al. (2025), and SCOPE Duong et al. (2025). It is worth noting that SCOPE also claims to require no external supervision. However, SCOPE tunes task-specific models by using gold reference labels as positive examples and synthesizes unfaithful negatives from a mixture of a fine-tuned model and an unconditional pre-trained language model. This process needs to be repeated for each type of downstream task. In contrast, SSFO is entirely label-free and pursues a more direct alignment strategy, yielding broadly task-agnostic faithfulness gains.
Full implementation details are provided in Appendix˜D.
@@ 4.1 Faithfulness Evaluation Results
Model Method Implement
@@
NQ-Swap Memo-Trap NQ-Open SQuAD Eli5 WikiQA
Span EM $\uparrow$ Span EM $\uparrow$ Span EM $\uparrow$ Span EM $\uparrow$ R-L F1 $\uparrow$ LFS $\uparrow$ R-L F1 $\uparrow$ LFS $\uparrow$
Llama-3-8B Instruct-Baseline \ \ 73.54% 73.60% 80.15% 88.20% 25.95% 59.80% 13.07% 75.31%
Decoding-Strategy CAD 74.83% 74.67% 69.30% 79.20% 21.40% 53.20% 15.02% 76.87%
DECORE 80.53% 74.40% 82.03% 84.90% 27.87% 68.90% 14.57% 78.41%
Post-Training ChatQA 67.70% 30.60% 76.80% 88.50% 27.13% 69.70% 13.83% 56.79%
Trust-Align 75.56% 70.95% 77.38% 50.90% 10.08% 55.10% 12.99% 76.19%
Context-DPO 82.76% 72.90% 82.86% 89.90% 27.19% 66.40% 11.00% 76.13%
SCOPE 76.72% 74.26% 80.38% 68.80% 22.41% 60.20% 15.69% 76.46%
SSFO 81.23% 76.28% 84.40% 89.00% 29.91% 71.40% 13.98% 75.72%
SSFO-$\lambda$ 82.81% 78.38% 85.69% 90.90% 31.48% 72.30% 15.53% 79.01%
Qwen2.5-7B Instruct-Baseline \ \ 79.35% 54.19% 82.29% 90.30% 23.11% 41.30% 15.33% 68.72%
Decoding-Strategy CAD 79.78% 63.10% 84.29% 85.90% 18.10% 49.80% 14.28% 26.75%
DECORE 81.93% 54.56% 83.76% 82.80% 26.83% 53.60% 14.49% 71.66%
Post-Training Trust-Align 79.69% 53.71% 77.93% 80.30% 15.67% 50.70% 16.30% 73.84%
Context-DPO 82.13% 55.34% 83.13% 91.80% 23.81% 49.30% 15.23% 72.84%
SCOPE 79.75% 44.90% 87.98% 78.50% 36.26% 60.20% 16.18% 51.03%
SSFO 84.18% 57.66% 83.88% 92.00% 24.49% 54.60% 15.96% 79.01%
SSFO-$\lambda$ 84.88% 60.77% 84.48% 93.30% 23.96% 62.80% 15.60% 74.07%
Mistral-7B Instruct-Baseline \ \ 67.76% 34.34% 79.13% 84.80% 23.38% 52.10% 18.34% 59.67%
Decoding-Strategy CAD 75.26% 22.57% 80.75% 89.00% 24.57% 48.30% 18.19% 32.92%
DECORE 78.17% 30.68% 86.52% 85.30% 25.24% 62.50% 22.07% 67.13%
Post-Training Context-DPO 79.62% 33.20% 80.68% 86.50% 24.55% 66.30% 15.29% 63.20%
SCOPE 49.58% 15.87% 64.71% 54.00% 27.06% 63.40% 15.47% 60.29%
SSFO 86.66% 37.22% 87.53% 89.00% 30.43% 80.60% 16.67% 63.79%
SSFO-$\lambda$ 85.48% 46.91% 90.32% 88.50% 33.58% 88.10% 19.64% 69.96%
*We present the results for varying model sizes in Table˜8.
@@
SSFO and SSFO-$\lambda$ deliver strong faithfulness across multiple datasets and models, as shown in Table˜1: Both variants of Self-Supervised Direct Preference Optimization (SSFO and SSFO-$\lambda$) substantially improve contextual faithfulness over the instruct baseline. For example, SSFO markedly improves Robustness, raising NQ-Swap Span EM scores by 18.9% for Mistral, and boosts Response Quality on NQ-Open for Llama from 80.15% to 84.40%. SSFO-$\lambda$ further increases scores on Llama and achieves top scores across all metrics for this model. Overall, SSFO-$\lambda$ attains state-of-the-art performance compared to both post-training-based baselines and decoding strategy-based baselines. These results indicate that explicitly encouraging benign likelihood displacement not only enhances faithfulness but also improves overall response quality.
@@
Table 2: A case study on faithfulness improvement. Comparing the instruct baseline, Context‐DPO, and SSFO on equivalent English–Spanish (ES) query–answer pairs. This example from XQuAD Artetxe et al. (2020) demonstrates that SSFO improves faithfulness over the instruct baseline and also preserves faithfulness cross-lingually. Blue texts highlight context-grounded information; Red texts mark unfaithful content.
@@
Instruct Model (EN) Cultural Imperialism is often referred to as cultural hegemony or cultural dominance. Hallucination
Context‑DPO (EN) Cultural Imperialism is often referred to as the spread of a dominant culture’s influence. This is also known as the use of soft power.                               Partially Context-Faithful
@@
[…]
Crose-language Response Quality Instruction Following
XQuAD(ES) DuReader(CN) FollowBench
Model Method Implement
Training Data
Required
Span EM $\uparrow$ Span EM $\uparrow$ CSL $\uparrow$
Instruct-Baseline \ \ 78.60% 78.80% 2.54
Decoding-Stratagy CAD \ 70.34% 76.57% 0.92
DECORE \ 81.87% 79.89% 2.46
ChatQA $\sim$30k 77.98% 72.05% 1.04
Trust-Align $\sim$15k 20.17% 8.12% 0.12
Context-DPO $\sim$5k 83.03% 84.40% 2.46
SCOPE $\sim$5k 69.70% 73.90% 0.16
SSFO $\sim$800 83.10% 84.90%   2.70
Llama-3-8B Post-Training SSFO-$\lambda$ $\sim$800 84.12% 83.56%   2.50
Instruct-Baseline \ \ 78.90% 81.50% 2.68
Decoding-Stratagy CAD \ 71.85% 76.71% 1.22
DECORE \ 80.08% 76.57% 2.56
Trust-Align $\sim$15k 75.21% 73.61% 0.58
Context-DPO $\sim$5k 79.92% 82.78% 2.64
SCOPE $\sim$5k 84.96% 89.27% 0.42
SSFO $\sim$800 79.83% 83.27%   2.70
Qwen2.5-7B Post-Training SSFO-$\lambda$ $\sim$800 81.76% 87.72%   2.62
@@ 4.2 Generalization Across Tasks and Languages
SSFO enhances multi-language faithfulness. We evaluate the generalization ability of SSFO in Table˜3, and results show it can improve cross-lingual faithfulness using only an English-based training set. For instance, on Llama, SSFO increases Span EM scores by 6.10% on DuReader (Chinese) and 5.52% on XQuAD (Spanish) compared to the instruct baseline. In contrast, heavily supervised methods like ChatQA Liu et al. (2025) and Trust-Align Song et al. (2025) exhibit decreased performance on these non-English QA datasets. This shows that by training the model to prioritize context knowledge over parametric knowledge, it learns a principle of contextual adherence that can transfer across languages.
SSFO minimally impacts instruction following capability. Requiring only a few hundred self-supervised data examples, SSFO largely preserves, and even slightly enhances, the model’s instruction following capabilities. The CSL scores on FollowBench indicate that models fine-tuned with SSFO retain comparable general instruction-following ability to the original base instruction models. We provide several cases that SSFO retains strong general instruction following ability under context-based scenarios in Table˜4. In contrast, other post-training approaches, such as Trust-Align Song et al. (2025), improve faithfulness at the cost of degrading general generative abilities (e.g., CSL score decreases from 2.54 to 0.12 on LLaMA-3-8B-Instruct).
Table 4: Case study from FollowBench Jiang et al. (2024). SSFO retains strong instruction-following capabilities under a context-based, composite NLP task. In contrast, the heavily post-trained Trust-Align model Song et al. (2025) fails, demonstrating a loss of general capabilities.
Query Analyze the provided text to pinpoint a category and the sentiment (positive, neutral, or negative) it emanates. Additionally, perform named entity recognition to emphasize notable entities and also identify the core topic discussed. Select the category from this array: company, educational institution, artist, athlete, office holder, means of transportation, building, natural place, village, animal, plant, album, film, or written work.
@@
Trust Align I apologize, but I couldn’t find an answer to your question in the search results. General Ability Lost
@@ 4.3 Data Efficiency Analysis
Figure 5: Data efficiency study: SSFO requires about 60% of data (400–500 examples) to achieve 85% of the total performance gain over the instruct baseline.
To measure how many self-supervised preference examples SSFO actually needs, we subsample the training dataset in 10% increments. As shown in Fig.˜5, we evaluate the average performance gain (an average improvement over the base instructed model). SSFO models cross the 85% performance threshold by approximately 50–60 % of the data (400–500 examples). We attribute this efficiency stems from using self-generated data, which avoids the stylistic distribution mismatch often caused by external data from human annotators or superior LLM models. Since the training data inherently matches the model’s native response style, optimization can focus on improving faithfulness. We compare the training examples from SSFO with other post-training methods in Table˜6.
@@ 5 Conclusion
This work addressed the critical challenge of faithfulness hallucination in RAG systems, where existing methods often introduce significant computational overhead or rely on costly external supervision. We introduced SSFO, an efficient self-supervised alignment approach that leverages the model’s own outputs to build preference pairs by comparing responses generated with retrieved context to responses based only on parametric knowledge. Our analysis shows the alignment proceeds through a benign form of likelihood displacement, which shifts probability mass from parametric-based tokens to context-aligned ones. Motivated by this finding, we proposed SSFO-$\lambda$, a variant that amplifies this beneficial displacement and further enhances faithfulness. Our experiments across diverse benchmarks show that SSFO and SSFO-$\lambda$ significantly enhance model faithfulness and robustness against parametric knowledge, achieving state-of-the-art performance compared to existing methods. Furthermore, SSFO exhibits strong generalization capabilities, improving faithfulness even in cross-lingual settings using only English training data, while preserving the model’s general instruction-following abilities.
Appendix A Appendix
Appendix B Related Work
B.1 Faithfulness Hallucination of Large Language Models
Hallucination in LLMs can be generally categorized into two types: factuality hallucination, where generated content deviates from established world knowledge (e.g., claiming “Mars has oceans”), and faithfulness hallucination, where the generated response is inconsistent with the provided context(e.g., misrepresenting a source document’s information). Huang et al.